# Repetition Threshold for Binary Automatic Sequences

### v2

@@ 1 Introduction
Let $x$ be a finite or infinite word, and $\alpha>1$ be real number. If $x$ has no finite factors of exponent $\geq\alpha$, we say it is $\alpha$-free. If it has no finite factors of exponent $>\alpha$, we say it is $\alpha^{+}$-free. We define the critical exponent of $\bf x$, written $\operatorname{ce}({\bf x})$, to be the supremum, over all finite nonempty factors $w$ of $\bf x$, of $\exp(w)$.
Let $S$ be a nonempty set of infinite words. We define the repetition threshold $\operatorname{rt}(S)$ to be the infimum, over all elements $\bf x$ of $S$, of $\operatorname{ce}({\bf x})$. Determining the repetition threshold of a class of words is an intriguing and often difficult problem. In its most famous incarnation, we take $S$ to be the set of all infinite words over a $k$-letter alphabet $\Sigma_{k}=\{0,1,\ldots,k-1\}$. In this case we define $\operatorname{RT}(k)=\operatorname{rt}(\Sigma_{k}^{\omega})$.
@@
Instead of studying the repetition threshold for all infinite words over a $k$-letter alphabet, we can also study various subsets of interest. For example, this was done for the case of Sturmian words [4], circular words [14], balanced words [21], sequences rich in palindromes [8], episturmian words [13], and other examples.
@@
We can also consider a more general notion, the generalized automatic sequences, where base-$k$ representation is replaced by other kinds of linear numeration systems, such as the Fibonacci numbers or Tribonacci numbers. For each numeration system, one can ask for its corresponding repetition threshold. In general, these problems seems quite difficult.
In this note we focus on the case of alphabet size $2$. Here we have a useful tool, a theorem that characterizes $\alpha$-free finite and infinite words [15]. There are three main results: Theorem 1, which exactly characterizes the repetition threshold for ${\cal A}_{k}$, $k\geq 2$, and Theorem 11, which does the same thing for the binary Fibonacci-automatic sequences, and Theorem 13 for Tribonacci-automatic sequences.
@@ 2 Results for $k$-automatic sequences
Let $i,j\geq 1$ be integers and $k\geq 2$. Then a sequence is $k^{i}$ automatic if and only if it is $k^{j}$ automatic.
@@
Let ${\bf x}$ be a generalized automatic sequence and let $w$ be a word. The lexicographically least sequence in the orbit closure of ${\bf x}$ beginning with $w$ is also a generalized automatic sequence for the same numeration system as ${\bf x}$.
@@
(a) If $k$ is a power of $2$, then by Lemma 4, a $2$-automatic sequence is also a $2^{i}$-automatic sequence for all $i\geq 1$. So it suffices to prove the result for $k=2$. But then $\bf t$ is a $2$-automatic sequence that avoids overlaps, and it is easy to see that no binary word of length $>3$ can avoid squares.
(b) Suppose $k$ is not a power of $2$. Suppose there is some ${\bf x}\in\operatorname{{\cal A}}_{k}$ with $\operatorname{ce}({\bf x})=\alpha<7/3$. Then, by Corollary 8, we know that $\bf x$ contains $\mu^{i}(0)$ as a factor for arbitrarily large $i$. Therefore it contains all sequences in the orbit closure of $\bf t$. By Lemma 9, the lexicographically smallest $7/3$-free sequence starting with $1$ is $\overline{\bf t}$, and hence the lexicographically smallest sequence in the orbit closure of ${\bf x}$ starting with $1$ is also $\overline{\bf t}$. It now follows from Lemma 10 that $\overline{\bf t}$ is $k$-automatic. Then, by a theorem of Cobham [5], we know (since $k$ and $2$ are multiplicatively independent) that $\overline{\bf t}$ must be ultimately periodic, a contradiction.
@@ 22 $0\rightarrow 0120102012030210201021$
Since the construction involving the morphism $h$ does not necessarily produce an automatic sequence with the smallest possible DFAO for a $k$-automatic $(7/3)^{+}$-free infinite binary word, it may be of interest to try to determine the number of states in a smallest DFAO. Lower bounds can be found by breadth-first search; upper bounds can be found by applying the construction involving $h$ or other means. The following table summarizes what we currently know. An asterisk denotes those entries where the exact value is known; intervals denote known lower and upper bounds. The automata can be constructed straightforwardly from the expression as the iteration of a fixed point starting with 00, followed by the specified coding.
@@
$k$ states automaton
$2$ $2$* $0\rightarrow 01$
@@
$\rightarrow 023$ 02302 (1)
$6$ $[6,27]$ apply $h$ to word generated by
morphism for $k=6$ in Table 1
$7$ $4$* $0\rightarrow 0120330$ (0)
@@ 3 Fibonacci- and Tribonacci-automatic sequences
In the Zeckendorf or Fibonacci numeration system [16, 27], we write every natural number $n$ uniquely as a sum of Fibonacci numbers $F_{i}$ for $i\geq 2$, subject to the condition that we never use both $F_{i}$ and $F_{i+1}$ in the sum. The representation can be viewed as a binary string, where we write $1$ if $F_{i}$ is included and 00 otherwise. A sequence $(a_{n})_{n\geq 0}$ is said to be Fibonacci-automatic if there is a DFAO that on input $n$ expressed in the Zeckendorf numeration system reaches a state with output $a_{n}$ [18, 10, 11]. We let $\operatorname{{\cal A}}_{F}$ denote the set of all binary Fibonacci-automatic sequences. The following is our second main result.
@@
Let $\bf x$ be Fibonacci-automatic. The possibility $\operatorname{ce}({\bf x})<7/3$ is ruled out by the same argument as in the proof of Lemma 10 (b) above using the morphic characterization of Fibonacci-automatic sequences in [23], and using Durand’s paper [12] instead of Cobham’s [5]. Namely, use Theorem 4.1 in [23] to obtain the Fibonacci-automatic sequence ${\bf x}$ as the pointwise image of a fixed point of a morphism $\varphi$ (as in Example 1 in [23, Section 5]). Check that the dominant eigenvalue of the incidence matrix of $\varphi$ is $\frac{1}{2}(1+\sqrt{5})$. Then use [12] since the numbers $2$ and $\frac{1}{2}(1+\sqrt{5})$ are multiplicatively independent.
@@
[…]
@@
$w_{n}$ =$=\begin{cases}w_{n-1}\overline{w_{n-2}}^{R},&\text{if $n\geq 2$ and $n\equiv 1,2\ \mbox{\rm(mod}\ 3\mbox{\rm)}$};\\[7.22743pt] w_{n-1}w_{n-2}^{R},&\text{if $n\geq 0\ \mbox{\rm(mod}\ 3\mbox{\rm)}$}.\end{cases}$
Clearly each $w_{i}$ is a prefix of all $w_{j}$ for $j\geq i$, so we can define $\bf w$ as the unique infinite word of which all the $w_{i}$ are prefixes. It is not difficult to show, with Walnut, that this word is generated by Fibonacci DFAO with $32$ states and is $(7/3)^{+}$-free. We omit the details. This means the size of the smallest such DFA is at most $32$. Breadth-first search shows it is at least $12$.
@@
It can be shown that the critical exponent of the sequence $\bf z$ is $1.875954364915\cdots$, the real zero of the polynomial $474X^{3}-24X^{2}+56X-$ 47.

### v3

@@ 1 Introduction
Let $x$ be a finite or infinite word, and $\alpha>1$ be real number. If $x$ has no finite factors of exponent $\geq\alpha$, we say it is $\alpha$-free. If it has no finite factors of exponent $\alpha$, we say it is $\alpha^{+}$-free. We define the critical exponent of $\bf x$, written $\operatorname{ce}({\bf x})$, to be the supremum, over all finite nonempty factors $w$ of $\bf x$, of $\exp(w)$.
Let $S$ be a nonempty set of infinite words (or infinite sequences; we use the two terms indifferently). We define the repetition threshold $\operatorname{rt}(S)$ to be the infimum, over all elements $\bf x$ of $S$, of $\operatorname{ce}({\bf x})$. Determining the repetition threshold of a class of words is an intriguing and often difficult problem. In its most famous incarnation, we take $S$ to be the set of all infinite words over a $k$-letter alphabet $\Sigma_{k}=\{0,1,\ldots,k-1\}$. In this case we define $\operatorname{RT}(k)=\operatorname{rt}(\Sigma_{k}^{\omega})$.
@@
Instead of studying the repetition threshold for all infinite words over a $k$-letter alphabet, we can also study various subsets of interest. For example, this was done for the case of Sturmian words [4], circular words [14], balanced words [21], sequences rich in palindromes [8], and episturmian words [13].
@@
We can also consider a more general notion, the generalized automatic sequences, where base $k$ is replaced by some linear numeration system. For each numeration system, one can ask for the corresponding repetition threshold for binary sequences. In general, these problems seems quite difficult. However, if the numeration system is addable—that is, the addition relation is computable by a finite automaton—then we have more tools at our disposal. Examples of this type of numeration system include the Zeckendorf system based on Fibonacci numbers, and an analogous system based on the Tribonacci numbers.
In this note we focus on the case of alphabet size $2$. Here we have a useful tool, a theorem that characterizes $\alpha$-free finite and infinite words [15]. There are three main results: Theorem 1, which exactly characterizes the repetition threshold for ${\cal A}_{k}$, $k\geq 2$; Theorem 11, which does the same thing for the binary Fibonacci-automatic sequences; and Theorem 13 for Tribonacci-automatic sequences.
@@ 2 Results for $k$-automatic sequences
Let $i,j\geq 1$ be integers and $k\geq 2$. Then a sequence is $k^{i}$-automatic if and only if it is $k^{j}$-automatic.
@@
We recall the notion of orbit closure of an infinite sequence $\bf x$: this is the set of all infinite sequences $\bf y$ such that every finite prefix of $\bf y$ appears somewhere in $\bf x$.
@@
Let $\bf x$} be a generalized automatic sequence in an addable numeration system and let $w$ be a finite word. The lexicographically least sequence in the orbit closure of $\bf x$} beginning with $w$ is also a generalized automatic sequence for the same numeration system as $\bf x$}.
@@
(a) If $k$ is a power of $2$, then by Lemma 4, a $2$-automatic sequence is also a $2^{i}$-automatic sequence for all $i\geq 1$. So it suffices to prove the result for $k=2$. But then $\bf t$ is a $2$-automatic sequence that avoids overlaps, and it is easy to see that no binary word of length >3>3 can avoid squares.
(b) Suppose $k$ is not a power of $2$. Suppose there is some ${\bf x}\in\operatorname{{\cal A}}_{k}$ with $\operatorname{ce}({\bf x})=\alpha<7/3$. Then, by Corollary 8, we know that $\bf x$ contains $\mu^{i}(0)$ as a factor for arbitrarily large $i$. Therefore the orbit closure of $\bf x$ contains all sequences in the orbit closure of $\bf t$. By Lemma 9, the lexicographically smallest $7/3$-free sequence starting with $1$ is $\overline{\bf t}$, and hence the lexicographically smallest sequence in the orbit closure of $\bf x$} starting with $1$ is also $\overline{\bf t}$. It now follows from Lemma 10 that $\overline{\bf t}$ is $k$-automatic. Then, by a theorem of Cobham [5], we know (since $k$ and $2$ are multiplicatively independent) that $\overline{\bf t}$ must be ultimately periodic, a contradiction.
@@ 22 $0\rightarrow 0120102012030210201021$
Since the construction involving the morphism $h$ does not necessarily produce an automatic sequence with the smallest possible DFAO for a $k$-automatic $(7/3)^{+}$-free infinite binary word, it may be of interest to try to determine the number of states in a smallest DFAO. Lower bounds can be found by breadth-first search; upper bounds can be found by applying the construction involving $h$ or other means. Table 2 summarizes what we currently know. An asterisk denotes those entries where the exact value is known; intervals denote known lower and upper bounds. The automata can be constructed straightforwardly from the expression as the iteration of a fixed point starting with 0, followed by the specified coding.
@@
$k$ states $s$ automaton
22 22* $0\rightarrow 01$
@@
$3\rightarrow 02302$ (1)
66 $6\leq s\leq 27$ Apply $h$ to word generated by
morphism for $k=6$ in Table 1.
77 44* $0\rightarrow 0120330$ (0)
@@ 3 Fibonacci- and Tribonacci-automatic sequences
In the Zeckendorf or Fibonacci numeration system [16, 27], we write every natural number $n$ uniquely as a sum of Fibonacci numbers $F_{i}$ for $i\geq 2$, subject to the condition that we never use both $F_{i}$ and $F_{i+1}$ in the sum. The representation can be viewed as a binary word, where we write $1$ if $F_{i}$ is included and 0 otherwise. A sequence $(a_{n})_{n\geq 0}$ is said to be Fibonacci-automatic if there is a DFAO that on input $n$ expressed in the Zeckendorf numeration system reaches a state with output $a_{n}$ [18, 10, 11]. We let $\operatorname{{\cal A}}_{F}$ denote the set of all binary Fibonacci-automatic sequences. The following is our second main result.
@@
Let $\bf x$ be Fibonacci-automatic. The possibility $\operatorname{ce}({\bf x})<7/3$ is ruled out by the same argument as in the proof of Lemma 10 (b) above. Namely, the automaton gives $\bf x$ as the image, under a coding, of a fixed point of some morphism $\varphi$ (as in Example 1 in [23, Section 5]). As in the base-$k$ case, the dominant eigenvalue of the incidence matrix of $\varphi$ is $\alpha=\frac{1}{2}(1+\sqrt{5})$. An easy induction proves that $\alpha^{i}=L_{i}/2+F_{i}\sqrt{5}/2$, where $L_{i}$ is the $i$’th Lucas number, so $\alpha$ and $2$ are multiplicatively independent. Then a theorem of Durand [12] finishes this part of the proof.
@@
There is such a sequence ${\bf w}=01101001101100110100\cdots$ generated by a DFAO of $32$ states. Let $x^{R}$ denote the reversal of the finite word $x$. We can construct $\bf w$ iteratively as follows: define
@@
$w_{n}$ $=\begin{cases}w_{n-1}\overline{w_{n-2}}^{R},&\text{if $n\geq 2$ and $n\equiv 1,2\ \mbox{\rm(mod}\ 3\mbox{\rm)}$};\\[7.22743pt] w_{n-1}w_{n-2}^{R},&\text{if $n\geq 0\ \mbox{\rm(mod}\ 3\mbox{\rm)}$}.\end{cases}$
Clearly each $w_{i}$ is a prefix of all $w_{j}$ for $j\geq i$, so we can define $\bf w$ as the unique infinite word of which all the $w_{i}$ are prefixes. It is not difficult to show, with Walnut, that this word is generated by a Fibonacci DFAO with $32$ states and is $(7/3)^{+}$-free. We omit the details. This means the size of the smallest such DFA is at most $32$. Breadth-first search shows it is at least $12$.
@@
It can be shown that the critical exponent of the sequence $\bf z$ is $1.875954364915\cdots$, the real zero of the polynomial $4X^{3}-24X^{2}+56X-47$.
@@ 4 Going further
Acknowledgments
We gratefully acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Nous remercions le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) de son soutien. The research of NR was supported by NSERC, grants RGPIN-2019-04111 and RGPIN-2025-04076. The research of JS was supported by NSERC, grant RGPIN-2024-03725.